OpenStudy (berrymox):

Show the following limit: lim x-> 0 (e^x -1) -------- x

OpenStudy (legomyego180):

do you know l'hopitals

OpenStudy (berrymox):

I can't use that, but yes

OpenStudy (berrymox):

I'm supposed to get e^0

OpenStudy (legomyego180):

what is e^0?

OpenStudy (berrymox):

1, but I don't know how to get it :P

OpenStudy (photon336):

\[e^{0} = 1\]

OpenStudy (legomyego180):

hm, but you cant use l'hopitals yet though

OpenStudy (berrymox):

Using classic algebra and limit laws.

OpenStudy (legomyego180):

This is indeterminate. There might be a way to rewrite using logarithmic properties but I've forgotten. If I were solving this I would take the derivative of both the numerator and denominator and plug in.

OpenStudy (photon336):

I have a weird idea maybe if we tried rationalizing the numerator @legomyego180 not sure if this would work \[\frac{ e^{x}-1 }{ x }*(\frac{ e^{x}+1 }{ e^{x}+1 }) = \frac{ e^{2x}-1 }{ x*e^{x}-1 }\]

OpenStudy (berrymox):

I went down that route, but it got too complicated. I might have done something wrong.

OpenStudy (legomyego180):

Actually, I dont think this is indeterminate. I think this DNE. @Photon336 thats certainly not wrong

OpenStudy (berrymox):

@Photon336 where do I go from here?

OpenStudy (photon336):

maybe insert x = 0 evaluate it and see what you get

OpenStudy (berrymox):

ok.

OpenStudy (legomyego180):

is this \[e^x-1 \] or e ^{x-1}

OpenStudy (berrymox):

I thought that it would just be 0/0, but I didn't look at it hard enough.

OpenStudy (legomyego180):

I dont think so.

OpenStudy (legomyego180):

I've never seen a conjugate used with e before but it works

OpenStudy (photon336):

@legomyego180 have you heard of wolfram alpha? You can check the result on there too

OpenStudy (legomyego180):

yup, I have pro if you want me to post the steps

OpenStudy (photon336):

yeah that would be great

OpenStudy (legomyego180):

Used L'hopitals but got the same answer

OpenStudy (berrymox):

e^(2*0) -1 = e^0 -1 = 1- 1 = 0

OpenStudy (photon336):

yeah, wow that's interesting

OpenStudy (berrymox):

How did I get -1 before? @Photon336

OpenStudy (photon336):

I ended up getting 1

OpenStudy (berrymox):

The final answer is 1, but I think I'm going crazy with simple exponent rules >_> for thinking that was -1. How do I prove the answer is 1 if your method doesn't work? (Unless it does, and I'm going crazy)

OpenStudy (berrymox):

*thinking numb. is -1

OpenStudy (photon336):

see what I did above

OpenStudy (berrymox):

I still don't understand, I'm getting 0/-1 = 0

OpenStudy (berrymox):

@Photon336

OpenStudy (photon336):

check this out. I found this alternative explanation online. e^x = 1 + x + x^2/2! + x^3/3! + .... => e^x - 1= x + x^2/2! + x^3/3! + .... => (e^x - 1)/x = 1 + x/2! + x^2/3! + ... => lim (x → 0) (e^x - 1) / x = lim (x → 0) [1 + x/2! + x^2/3! + ...] = 1.

OpenStudy (photon336):

@Berrymox so my guess here is that as x->0 and get's really really small, all those terms I guess converge to 1

OpenStudy (berrymox):

I feel bad for you typing all that out, but I can't use that method either.

OpenStudy (berrymox):

*At least, not in my homework

OpenStudy (photon336):

@Berrymox what methods did you learn in class?

OpenStudy (berrymox):

just the methods you were doing before with rationalizing and stuff. Basic algebra.

OpenStudy (photon336):

well, we gotta figure out something

OpenStudy (berrymox):

I'll just leave it as a graph and take the point off then

OpenStudy (zarkon):

use the definition of the derivative assuming you know the derivative of \(e^x\) \[\lim_{x\to0}\frac{e^x-1}{x}=\lim_{x\to0}\frac{e^x-e^0}{x-0}=\left.\frac{d}{dx}e^{x}\right|_{x=0}=e^0=1\]

OpenStudy (photon336):

fascinating

OpenStudy (zarkon):

\[\lim_{x\to a}\frac{f(x)-f(a)}{x-a}=f'(a)\] here \(f(x)=e^x\) and \(a=0\)